# Humanoid Policy $\sim$ Human Policy

### v2

@@ IV Experiments
TABLE II: Success rate of autonomous skill execution. Co-training with human data (H. Data) significantly improves the Out-Of-Distribution (O.O.D.) performance with nearly 100% relative improvement on all tasks. We also ablate the design choice of using different normalizations (D. Norm) for different embodiments.
This section investigates how cross-embodiment learning with egocentric task-oriented human demonstrations improves the robustness and generalizability of acquired policies, which highlights the potential of PH2D as a scalable data source. We also provide qualitative examples of manipulation tasks and ablate design choices.
@@
We run our experiments on Unitree H1 bimanual humanoid robots equipped with a pair of 6-DOF Inspire dexterous hands. We focus on manipulation and do not actuate the lower body of the robot. Instead, we introduce a 2 DoF actuated neck similar to the design in [9], allowing the robot to visually focus on the objects it manipulates, similar to how humans do. This configuration results in a total of 28 degrees of freedom. The robot setup is illustrated in Fig. 3 and Fig. 4. Note that we rely solely on head cameras, without using wrist cameras, to study the manipulation capabilities based purely on egocentric vision.
@@
We implement policy architecture by adopting the implementation from OpenTV [9]. In particular, we replace the trainable ImageNet-pretrained ResNet-18 backbone with frozen DinoV2 [43] ViT-S for its improved robustness against lighting and texture changes. Unless specifically noted, we use different means and standard deviations to normalize the state-action space. We implement two variants:
  • ACT: baseline implementation using the architecture described above, trained using only robot data. Robot states are represented as joint positions.
  • HAT: same architecture as ACT, but the state encoder operates in the unified state-action space. Unless otherwise stated, HAT is co-trained on robot and human data.
@@
We collect robot and human demonstrations in different object sets. Since human demonstrations are easier to collect, the settings in human demonstrations are generally more diverse, which include background, object types, object positions, and the relative position of the human to the table. We experimented with four different dexterous manipulation tasks and investigated in-distribution and out-of-distribution setups. The in-distribution (I.D.) setting tests the learned skills with backgrounds and object arrangements approximately similar to the training demonstrations presented in the real-robot data. In the Human-Distribution (H.D.) setting, we evaluate on-scene setups included in the human demonstrations but not robot demonstrations. In the Out-Of-Distribution (O.O.D.) setting, we test generalizability and robustness by introducing novel setups that were not present in any training demonstrations. Fig. 4 visualizes different manipulation tasks and how we define out-of-distribution settings for each task.
With the setup above, we aim to answer important research questions as follows:
  • Does co-training with human data improve I.D. performance?
  • Does co-training with human data improve O.O.D. generalization?
  • How efficient is human demonstration collection compared to teleoperation [9]?
  • How much does each design choice contribute?
IV-A Overall Evaluation
@@
TABLE III: Background Generalization: In the cup passing task, we evaluate the passing performance by recording the number of failures or retries needed to complete 20 cup-passing trials.
Human data has minor effects on I.D. testing. From Tab. II, we can see that I.D. performance with or without co-training with human data gives similar results. In the I.D. setting, we attempt to replicate the scene setups as training demonstrations, including both background, object types, and object placements. Thus, policies trained with only a small amount of robot data performed well in this setting. This finding is consistent with recent work [9, 11] that frozen visual foundation models [46, 43] improve robustness against certain perturbations such as lighting and similar textures.
Human data improves the O.O.D. settings. One common challenge in imitation learning is overfitting to only in-distribution task settings. Hence, it is crucial for a robot policy to generalize beyond the scene setups seen in a limited set of single-embodiment data. To demonstrate how co-training with human data reduces such overfitting, we introduce O.O.D. task setting to evaluate such generalization. From Tab. II, we can see that co-training drastically improves O.O.D. settings, achieving nearly 100% relative improvement in settings unseen by the robot data. In particular, we find that human data improves three types of generalization: background, object placement, and appearance. To isolate the effect of each variable, we evaluate different types of generalization in separate tasks. We provide in-depth analyses of each type of generalization below.
Method Bottle Box1 Box2 Can Ovr. Succ.
I.D. H.D. O.O.D. O.O.D.
ACT 8/10 5/10 1/10 1/10 16/40
HAT 8/10 7/10 1/10 4/10 21/40
TABLE IV: Object Appearance Generalization: In the horizontal grasping task, we evaluated the grasping performance by attempting to grasp each object 10 times and recorded the success rate.
IV-B Cross-embodiment Generalization
Human data improves background generalization. We chose to use the cup passing task to test background generalization. We prepared four different tablecloths as backgrounds, as shown in Fig. 4(a).
In terms of training data distribution, the teleoperation data for this task was collected exclusively on the paper background shown in Fig. 4(a), whereas the human data includes more than five different backgrounds. This diverse human dataset significantly enhances the generalization ability of the co-trained HAT policy. As shown in Tab. III. , HAT consistently outperforms across all four backgrounds, demonstrating robustness to background variations. In addition, the overall success rate increases by nearly 50% compared to training without human data, highlighting the advantage of utilizing diverse human demonstrations.
Since grasping is a relatively simple task, our adjusted policy demonstrates strong learning capabilities even with only 50 teleoperation data samples. The policy can successfully grasp most bottles despite the limited training set. To better highlight the impact of human data, we selected more challenging objects for evaluation. As shown in Tab. IV, human data significantly enhances the policy’s ability to grasp these more difficult objects.
Notably, box_1 appears in the human data, while box_2 does not. Despite this, we observe that co-training with human data still improves overall performance, even on box_2, though its success rate does not increase. This suggests that, beyond direct experience with specific objects, the human data helps the policy learn broader visual priors that enable more proactive and stable grasping behaviors. For box_2, while the success rate remains low—partially due to its low height and color similarity to the table—the co-trained HAT policy demonstrates fewer out-of-distribution (OOD) failures and more actively searches for graspable regions. The failures on box_2 are primarily due to execution errors in grasping rather than the inability to perceive or locate the object.
Method Bottle Box1 Box2 Can Ovr. Succ.
I.D. H.D. H.D. H.D.
Without whole-body 8/10 6/10 0/10 7/10 21/40
With whole-body 9/10 3/10 3/10 3/10 18/40
TABLE V: Ablation of how human whole-body movement in training demonstrations affects policy rollout. We collect the same number of demonstrations on the same set of objects for the grasping task with or without whole-body movement. Since the robot does not have a natural whole-body movement like humans, it negatively influences the manipulation success rate.
IV-C Ablation Study
Normalization of different embodiments. Tab. II suggests minor differences between using different normalizations of states and actions for humans and humanoids. We take a closer look in Fig. 5, where we investigate the impact of different normalization strategies in the vertical grasping (picking) task. Noticeably, the same normalization approach achieved the highest overall success rate, but the success distribution is biased towards the upper-right region of the grid.
Autonomous Whole-body Movement. In Tab. V, we justify the necessity to minimize body movement in human data collection. Humans tend to move their upper body unconsciously during manipulation (including shoulder and waist movement). However, existing humanoid robots have yet to reach such a level of dexterity. Thus, having these difficult-to-replicate actions in the human demonstrations leads to degraded performance. We hypothesize that such a necessity would be greatly reduced with the development of both whole-body locomotion methods and mechanical designs, but for the currently available platforms, we instruct operators to minimize body movement as much as possible in our dataset.
State-Action Design. In Tab. VI, we ablate the design choices of the proprioception state space and the speed of output actions. In particular, using the same set of robot and human data, we implement two baselines: 1) a unified state-action space, but does not interpolate (i.e., slow down) the human actions; and 2) a baseline that interpolates human actions but uses separate state representation for humanoid (joint positions) and humans (EEF representation). The policies exhibit different failure patterns during the rollout of these two baselines. Without interpolating human actions, the speed of the predicted actions fluctuates between fast (resembling humans) and slow (resembling teleoperation), which leads to instability. Without a unified state space, the policy implicitly learns to differentiate between different embodiments. Rather than developing a shared representation across embodiments, it appears to rely on a form of conditional switching—activating the corresponding embodiment-specific policy when observing a familiar state. While this approach results in high success rates for in-distribution evaluation, it provides limited benefits for cross-embodiment generalization, as the policy struggles to transfer knowledge effectively between different embodiments.
Task State Space Action Speed Success
1/10
Vertical Grasping 0/10
4/10
TABLE VI: Importance of unifying policy inputs and outputs. We report the number of successes of vertical grasping objects in the upper-left block as illustrated in Fig. 5. Baselines use joint positions as state input or do not interpolate human motions.
Efficiency of Data Collection. In Tab. VII, we compare task completion times across different setups, including standard human manipulation, human demonstrations performed while wearing a VR device, and robot teleoperation. This analysis highlights how task-oriented human demonstrations can be a scalable data source for cross-embodiment learning. Notably, wearing a VR device does not significantly impact human manipulation speed, as the completion time remains nearly the same as in standard human demonstrations.
Method Grasping (secs) Pouring (secs)
Human Demo 3.79$\pm$0.27 4.81$\pm$0.35
Human Demo with VR 4.09$\pm$0.30 4.90$\pm$0.26
Humanoid Demo (VR Teleop) 19.72$\pm$1.65 37.31$\pm$6.25
TABLE VII: Amortized mean and standard deviation of the time required to collect a single demonstration, including scene resets. The first row shows the time for regular human to complete corresponding tasks in real world. The second row represents our human data when wearing VR for data collection, demonstrating that egocentric human demonstrations provide a more scalable data source compared to robot teleoperation.
IV-D Few-Shot Transfer across Heterogenous Embodiments
Figure 6: Co-training Efficacy and Generalization across Objects. Performance of Humanoid B on co-training with Humanoid A and Human on vertical grasping across 4 objects. ID o1 is an object common to all embodiments. OOD o2 and OOD o3 are from Humanoid A demonstrations and Human demonstrations respectively. OOD o4 is out-of distribution for all embodiments.
To further evaluate our method’s cross-embodiment adaptability, we conducted few-shot generalization experiments on a distinct humanoid platform (Unitree H1_2, termed Humanoid B), contrasting it with our primary platform (modified Unitree H1, Humanoid A). Notably, Humanoid B’s demonstration data were collected in an entirely separate environment with only a single object, introducing both embodiment and environmental shifts.
Experiment 1: Cross-embodiment co-training efficacy Using only 20 demonstrations from Humanoid B, we trained 3 policies - respectively on data from (i) Humanoid B only, (ii) Humanoid B + Humanoid A (cross-embodiment), and (iii) Humanoid B + Humanoid A + Human (cross-embodiment and human priors). As shown in Fig. 6 co-training policies (ii) and (iii) substantially outperformed the Humanoid B-only baseline. Notably, co-training improved success rates not only on in-distribution objects (seen by Humanoid B) but also on novel objects never encountered by Humanoid B during training, underscoring the method’s ability to transfer latent task structure across embodiments.
Experiment 2: Scaling Demonstrations for Few-Shot Adaptation We further quantified the relationship between required for few-shot generalization. While holding Humanoid A (100 demos) and human (50 demos) datasets fixed, we incrementally increased Humanoid B’s demonstrations (Fig. 7). Co-training (Humanoid B + A + Human) consistently outperformed isolated training on Humanoid B across all data regimes. Crucially, even with as few as 5–10 demonstrations from Humanoid B, co-training achieved more than twice the successful rollouts on novel objects, whereas zero-shot transfer (isolated training) remained near chance (12%). This demonstrates that our method enables data-efficient adaptation: by leveraging shared representations from diverse embodiments, minimal target-domain demonstrations suffice to anchor robust policies.
These results highlight two key advantages of our approach: (1) the ability to unify heterogeneous demonstration sources (robots, humans) into a generalizable policy framework, and (2) the capacity to rapidly adapt to new embodiments with drastically reduced data requirements. Such capabilities are critical for scaling robot learning to real-world settings, where per-platform data collection is often prohibitively expensive.
V Conclusions
VI Limitation
Although we also collect language instructions in PH2D, due to our focus on investigating the embodiment gap between humans and humanoids, one limitation of the current version of the paper uses a relatively simple architecture for learning policy. In the near future, we plan to expand the policy learning process to train a large language-conditioned cross-embodiment policy to investigate generalization to novel language using human demonstrations. In addition, current evaluations are done on robots equipped with dexterous hands, which are more aligned with humans. One future direction is to validate whether PH2D also facilitates cross-embodiment learning of other forms of robots, such as those with parallel grippers.

### v3

@@ 4 Experiments
Type of Generalization Background Texture Obj. Placement
Table 2: Success rate of autonomous skill execution. Co-training with human data (H. Data) significantly improves the Out-Of-Distribution (O.O.D.) performance with nearly 100% relative improvement on all tasks on Humanoid A. We also ablate the design choice of using different normalizations (D. Norm) for different embodiments. We designate each task setting to investigate a single type of generalization. Detailed analysis of each type of generalization is presented in Sec. C.
Figure 4: Hardware Illustration. Most robot data attributes to Humanoid A, a Unitree H1 robot. Humanoid B, a Unitree H1-2 robot with different arm motor configurations, is used to evaluate few-shot cross-humanoid transfer. Detailed comparisons in Sec. D
@@
We run our experiments on two humanoid robots (Humanoid A and Humanoid B shown in Fig. 4) equipped with 6-DOF Inspire dexterous hands. Humanoid A is a Unitree H1 robot and Humanoid B is a Unitree H1_2 robot with different arm configurations. Similar to humans, both robots (1) are equipped with actuated necks [9] to get make use of egocentric views and (2) do not have wrist cameras. Unless otherwise noted, most humanoid data collection is done with Humanoid A. We use Humanoid B mainly for testing cross-humanoid generalization.
@@
We implement policy architecture by adopting an transformer-based architecture predicting future action chunks [5]. We use a frozen DinoV2 ViT-S [67] as the visual backbone. We implement two variants: (1) ACT: baseline implementation using the Action Chunk Transformer [5], trained using only robot data. Robot states are represented as joint positions. (2) HAT: same architecture as ACT, but the state encoder operates in the unified state-action space. Unless otherwise stated, HAT is co-trained on robot and human data. A checkpoint is trained for each task with approximately 250-400 robot demonstrations.
@@
We collect robot and human demonstrations in different object sets. Since human demonstrations are easier to collect, the settings in human demonstrations are generally more diverse, which include background, object types, object positions, and the relative position of the human to the table.
(a) Performance of Humanoid B co-trained with PH2D on horizontal grasping. o1 is seen by Humanoid B. o2 and o3 seen in human data. o4 is unseen in all data.
(b) Co-training consistently outperforms isolated training as Humanoid B demonstrations increase, achieving good success rates even in low-data regimes.
@@
We experimented with four different dexterous manipulation tasks and investigated in-distribution and out-of-distribution setups. The in-distribution (I.D.) setting tests the learned skills with backgrounds and object arrangements approximately similar to the training demonstrations presented in the real-robot data. In the Out-Of-Distribution (O.O.D.) setting, we test generalizability and robustness by introducing novel setups that were presented in human data but not in robot data. Fig. 7 visualizes different manipulation tasks and how we define out-of-distribution settings for each task.
4.1 Main Evaluation
Human data has minor effects on I.D. testing. From Tab. 2, we can see that I.D. performance with or without co-training with human data gives similar results. In the I.D. setting, we closely match the scene setups as training demonstrations, including both background, object types, and object placements. Thus, policies trained with only a small amount of Humanoid A data performed well in this setting. This finding is consistent with recent work [9, 7] that frozen visual foundation models [17, 67] improve robustness against certain external perturbations such as lighting.
Human data improves the O.O.D. settings with many generalizations. One common challenge in imitation learning is overfitting to only in-distribution task settings. Hence, it is crucial for a robot policy to generalize beyond the scene setups seen in a limited set of single-embodiment data. To demonstrate how co-training with human data reduces such overfitting, we introduce O.O.D. task settings to evaluate such generalization. From Tab. 2, we can see that co-training drastically improves O.O.D. settings, achieving nearly 100% relative improvement in settings unseen by the robot data. In particular, we find that human data improves three types of generalization: background, object placement, and appearance. To isolate the effect of each variable, each task focuses on a specific type of generalization as listed in Tab. 2, with in-depth analyses in Sec. C.
4.2 Few-Shot Transfer across Heterogenous Embodiments
We conducted few-shot generalization experiments on a distinct humanoid platform (Humanoid B), contrasting it with our primary platform, Humanoid A. Notably, Humanoid B’s demonstration data were collected in an entirely separate environment, introducing both embodiment and environmental shifts. We highlight two key advantages of our approach: (1) the ability to unify heterogeneous human-centric data sources (humanoids and humans) into a generalizable policy framework, and (2) the capacity to rapidly adapt to new embodiments with drastically reduced data requirements.
Figure 6: Human data has better sampling efficiency. Per-grid vertical grasping successes out of 10 trials with models trained with robot-only data and mixed data. Red boxes indicate where training data is collected.
Task State Space Action Speed Success
1/10
Vertical Grasping 0/10
4/10
Table 3: Importance of unifying policy inputs and outputs. We report the number of successes of vertical grasping objects in the upper-left block as illustrated in Fig. 8. Baselines use joint positions as state input or do not interpolate human motions.
Experiment 1: Cross-embodiment co-training efficacy Using only 20 demonstrations from Humanoid B, we trained 3 policies - respectively on data from (i) Humanoid B only, (ii) Humanoid B + Humanoid A (cross-embodiment), and (iii) Humanoid B + Humanoid A + Human (cross-embodiment and human priors). As shown in Fig. 5(a), co-trained policies (ii) and (iii) substantially outperformed the Humanoid B-only baselines on all task settings, underscoring the method’s ability to transfer latent task structure across embodiments.
Experiment 2: Scaling Demonstrations for Few-Shot Adaptation We further quantified the relationship between required for few-shot generalization. We hold Humanoid A and human datasets fixed for the horizontal grasping task and ablate number of demonstrations required for Humanoid B in Fig. 5. Co-training (Humanoid B + A + Human) consistently outperformed isolated training on Humanoid B across all settings, especially in the few-data regime.
4.3 Ablation Study
Sampling Efficiency of Human and Humanoid Data. Conceptually, collecting human data is less expensive, not just because it can be done faster, but also because it can be done in in-the-wild scenes; reduces setup cost before every data collection; and avoids the hardware cost to equip every operator with robots.
We perform additional experiments to show that even in the lab setting, human data can have better sampling efficiency in unit time. In particular, we provide a small-scale experiment on the vertical grasping task. Allocating 20 minutes for two settings, we collected (1) 60 Humanoid A demonstrations, (2) 30 Humanoid A demonstrations, and 120 human demonstrations. To avoid conflating diversity and data size, the object placements in all demonstrations are evenly distributed at the bottom 6 cells. The results are given in Fig. 6. The policy trained with mixed robot and human data performs significantly better, which validates the sampling efficiency of human data over robot data. Each cell represents a 10cm × 10cm region where the robot attempts to pick up a box.
State-Action Design. In Tab. 3, we ablate the design choices of the proprioception state space and the speed of output actions. In particular, using the same set of robot and human data, we implement two baselines: 1) a unified state-action space, but does not interpolate (i.e., slow down) the human actions; and 2) a baseline that interpolates human actions but uses separate state representation for humanoid (joint positions) and humans (EEF representation). The policies exhibit different failure patterns during the rollout of these two baselines. Without interpolating human actions, the speed of the predicted actions fluctuates between fast (resembling humans) and slow (resembling teleoperation), which leads to instability. Without a unified state space, the policy is given a ‘shortcut’ to distinguish between embodiments, which leads to on-par in-distribution performance and significantly worse OOD performance.
More Ablation Study. Due to space limit, please refer to the appendix and the supplementary for more qualitative visualization and quantitative ablation studies.
5 Conclusions
6 Limitations
Although we also collect language instructions in PH2D, due to our focus on investigating the embodiment gap between humans and humanoids, one limitation of the current version of the paper uses a relatively simple architecture for learning policy. In the near future, we plan to expand the policy learning process to train a large language-conditioned cross-embodiment policy to investigate generalization to novel language using human demonstrations. The collection of human data relies on off-the-shelf VR hardwares and their hand tracking SDKs. Since these SDKs were trained mostly for VR applications, hand keypoint tracking can fail for certain motions with heavy occlusion. In addition, though the proposed method conceptually extends to more robot morphologies, current evaluations are done on robots equipped with dexterous hands.
7 Acknowledgment
This work was supported, in part, by NSF CAREER Award IIS-2240014, NSF CCF-2112665 (TILOS), and gifts from Amazon, Meta and Apple.
Appendix A More Ablation Study - Data Collection
Method Bottle Box1 Box2 Can Ovr. Succ.
I.D. H.D. H.D. H.D.
Without whole-body 8/10 6/10 0/10 7/10 21/40
With whole-body 9/10 3/10 3/10 3/10 18/40
Table 4: Ablation of how human whole-body movement in training demonstrations affects policy rollout. We collect the same number of demonstrations on the same set of objects for the grasping task with or without whole-body movement. Since the robot does not have a natural whole-body movement like humans, it negatively influences the manipulation success rate.
Autonomous Whole-body Movement. In Tab. 4, we justify the necessity to minimize body movement in human data collection. Humans tend to move their upper body unconsciously during manipulation (including shoulder and waist movement). However, existing humanoid robots have yet to reach such a level of dexterity. Thus, having these difficult-to-replicate actions in the human demonstrations leads to degraded performance. We hypothesize that such a necessity would be greatly reduced with the development of both whole-body locomotion methods and mechanical designs, but for the currently available platforms, we instruct operators to minimize body movement as much as possible in our dataset.
Efficiency of Data Collection. In Tab. 5, we compare task completion times across different setups, including standard human manipulation, human demonstrations performed while wearing a VR device, and robot teleoperation. This analysis highlights how task-oriented human demonstrations can be a scalable data source for cross-embodiment learning. Notably, wearing a VR device does not significantly impact human manipulation speed, as the completion time remains nearly the same as in standard human demonstrations.
Method Grasping (secs) Pouring (secs)
[…]
Table 5: Amortized mean and standard deviation of the time required to collect a single demonstration, including scene resets. The first row shows the time for regular human to complete corresponding tasks in real world. The second row represents our human data when wearing VR for data collection, demonstrating that egocentric human demonstrations provide a more scalable data source compared to robot teleoperation.
Appendix B Normalization of different embodiments.
Tab. 2 suggests minor differences between using different normalization coefficients for the states and actions vectors of humans and humanoids. We take a closer look in Fig. 8, where we investigate the impact of different normalization strategies in the vertical grasping (picking) task. Noticeably, the same normalization approach achieved the highest overall success rate, but the success distribution is biased towards the upper-right region of the grid.
Appendix C In-Depth Analysis of Different Types of Generalization
Human data improves background generalization. We chose to use the cup passing task to test background generalization. We prepared four different tablecloths as backgrounds, as shown in Fig. 7(a). In terms of training data distribution, the teleoperation data for this task was collected exclusively on the paper background shown in Fig. 7(a), whereas the human data includes more than five different backgrounds. This diverse human dataset significantly enhances the generalization ability of the co-trained HAT policy. As shown in Tab. 7. , HAT consistently outperforms across all four backgrounds, demonstrating robustness to background variations. In addition, the overall success rate increases by nearly 50% compared to training without human data, highlighting the advantage of utilizing diverse human demonstrations.
Since grasping is a relatively simple task, our adjusted policy demonstrates strong learning capabilities even with only 50 teleoperation data samples. The policy can successfully grasp most bottles despite the limited training set. To better highlight the impact of human data, we selected more challenging objects for evaluation. As shown in Tab. 6, human data significantly enhances the policy’s ability to grasp these more difficult objects.
Notably, box_1 appears in the human data, while box_2 does not. Despite this, we observe that co-training with human data still improves overall performance, even on box_2, though its success rate does not increase. This suggests that, beyond direct experience with specific objects, the human data helps the policy learn broader visual priors that enable more proactive and stable grasping behaviors. For box_2, while the success rate remains low—partially due to its low height and color similarity to the table—the co-trained HAT policy demonstrates fewer out-of-distribution (OOD) failures and more actively searches for graspable regions. The failures on box_2 are primarily due to unstable grasping and the small box slipping from the hand, rather than the inability to perceive or locate the object.
Method Bottle Box1 Box2 Can Ovr. Succ.
I.D. H.D. O.O.D. O.O.D.
ACT 8/10 5/10 1/10 1/10 16/40
HAT 8/10 7/10 1/10 4/10 21/40
Table 6: Object Appearance Generalization: In the horizontal grasping task, we evaluated the grasping performance by attempting to grasp each object 10 times and recorded the success rate.
@@ Appendix C In-Depth Analysis of Different Types of Generalization
Table 7: Background Generalization: In the cup passing task, we evaluate the passing performance by recording the number of failures or retries needed to complete 20 cup-passing trials.
Appendix D In-Depth Comparison between Humanoid A and Humanoid B configurations
This section presents a detailed comparison of the two humanoid platforms, referred to as Humanoid A and Humanoid B, with a focus on joint structure and implications for manipulation capabilities. We restrict our analysis to the arm configurations, as other parts of the body were not exclusively explored in this work.
While morphologically similar, these two humanoids have drastically different arm configurations that create hurdles in direct policy transfer. Besides differences in motor technical specs such as torque and types of encoder (Humanoid B has absolute motor position encoders), they also have different mechanical limits. The range of motion (ROM) for the first four proximal joints—shoulder_pitch, shoulder_roll, shoulder_yaw, and elbow—differs across the two platforms. Humanoid B exhibits a consistently wider ROM, which allows a wider set of reachable configurations and increases the manipulability of the arm in constrained environments. Table 8 summarizes the ROM values for these shared joints.
A more significant architectural divergence is observed at the wrist. Humanoid A includes a single distal joint—wrist_roll—providing limited wrist articulation. This restricts end-effector dexterity and constrains in-hand manipulation strategies to a single rotational degree of freedom. In contrast, Humanoid B is equipped with a complete wrist mechanism composed of three independently actuated joints: wrist_pitch, wrist_roll, and wrist_yaw. These additional degrees of freedom allow for full orientation control of the end-effector, enabling tasks that require precise alignment, rotation, and fine adjustment of object poses.
Joint Humanoid A Humanoid B
shoulder_pitch $164^{\circ}$ to $164^{\circ}$ $180^{\circ}$ to $90^{\circ}$
shoulder_roll $19^{\circ}$ to $178^{\circ}$ $21^{\circ}$ to $194^{\circ}$
shoulder_yaw $74^{\circ}$ to $255^{\circ}$ $152^{\circ}$ to $172^{\circ}$
elbow $71^{\circ}$ to $150^{\circ}$ $54^{\circ}$ to $182^{\circ}$
wrist_roll $175^{\circ}$ to $175^{\circ}$ $172^{\circ}$ to $157^{\circ}$
Table 8: Joint Range of Motion Comparison between Humanoid A and B (in degrees)